OpenStudy (anonymous):

You have already invested $600 in a stock with an annual return of 11%. How much of an additional $1,200 should be invested at 12% and how much at 6% so that the total return on the entire $1,800 is 9%? (Round each answer to the nearest cent.)

OpenStudy (istim):

So, what have you done so far?

OpenStudy (rogue):

600 * 0.11 = $66 1800 * 0.09 = $162 The return from the 1200 should be 162 - 66 = $96 x + y = 1200 .12x + .6y = 162 Solve the system.

OpenStudy (rogue):

Can you do that?

OpenStudy (rogue):

oops, .12x + .6y = 96, not 162

OpenStudy (anonymous):

so from there i just solve for Y and that's my answer?

OpenStudy (rogue):

After you solve for y, use that to solve for x as well. y is the amount that is going to be invested at 6%, x is invested at 12%

OpenStudy (anonymous):

wait now im lost haha

OpenStudy (rogue):

x + y = 1200 .12x + .06y = 96 x = 1200 - y ------> .12x + 0.06y = 96 .12(1200 - y) + 0.06y = 96 144 - .12y + 0.06y = 96 48 = 0.06y y = 800 x + y = 1200 x + 800 = 1200 x = 400

OpenStudy (rogue):

Sorry, I made a mistake (twice :/) originally when I posted the system. I put 0.6y instead of 0.06y.